# Hyperbolic Dataset Distillation

### v1

@@ A.1 Tangent Space $T_{\mathbf{x}}\mathcal{L}$
In the Lorentz model, hyperbolic space $\mathcal{L}$ is realized as a sheet of the two-sheeted hyperboloid in $\mathbb{R}^{n+1}$ with Minkowski metric. For any point $\mathbf{x}=[x_{t};x_{s}]\in\mathcal{L}$, the tangent space captures all possible instantaneous directions at $\mathbf{x}$. It is defined by
@@ A.2 Exponential and Logarithm Maps
The exponential map pushes vectors in the tangent space onto the manifold, yielding a local Euclidean-like parametrization. Let $\kappa=\sqrt{-K}$. For $\mathbf{v}\in T_{\mathbf{x}}\mathcal{L}$, define
@@ A.3 Bijection between the Lorentz and Poincaré Ball Models
For many applications (especially visualization) it is convenient to switch to the Poincaré ball. Given a Lorentz point $\mathbf{x}=[x_{t};x_{s}]$, we map it to the unit ball $\|\mathbf{p}\|<1$ via
@@
One verifies that the reconstructed $\mathbf{x}$ satisfies $-x_{t}^{2}+\|x_{s}\|^{2}=1/K$ and $x_{t}>0$.
@@ Appendix B Centroid Convergence Toward the Origin
Since $\mathbb{L}_{K}^{n}$ has constant curvature $K<0$, each map $\mathbf{p}\mapsto\|\ Log_{\mathbf{p}}(\mathbf{p}_{i})\|^{2}$ is strictly convex along geodesics, ensuring a single global minimizer. We choose the origin $P_{0}$ to be the unique fixed point of a maximal compact subgroup of $\ Isom(\mathbb{L}_{K}^{n})$, whose stabilizer is isomorphic to $\mathrm{O}(n)$. A comparison-theorem argument then shows
@@
forcing the solution of $\sum_{i}\ Log_{\mathbf{p}}(\mathbf{p}_{i})=0$ to lie radially closer to $p_{0}$ than the Euclidean centroid. Moreover, as $\lvert K\rvert$ increases, the lower bound on the second-derivative of $t\mapsto\|\ Log_{\gamma(t)}(\mathbf{p}_{i})\|^{2}$ along any geodesic $\gamma$ grows, making this radial bias toward $p_{0}$ even more pronounced. This results in the centroids of both the original dataset and the synthetic dataset being biased towards $p_{0}$, while the distance between them is relatively small.
@@ Appendix C Hierarchical Weight
:=$\mathcal{L}:=\bigl{\{}x\in\mathbb{R}^{n+1}\mid\langle x,x\rangle_{\mathcal{L}}=-R^{2},\;x_{0}>0\bigr{\}},$ (31)
where $\langle\cdot,\cdot\rangle_{L}$ denotes the Minkowski inner product of signature (++)(-$\cdots$+)( - + ⋯ + ). The geodesic distance between two points $p,q\in\mathbb{H}_{K}^{n}$ is given by
@@ Appendix D Gradient Contributions in the Lorentz Model of Hyperbolic Space
[…]
Thus points very close to $x$ exert almost the maximal pull of magnitude 1.
@@ Appendix E Benchmark Datasets
FashionMNIST is a drop-in replacement for the classic MNIST dataset, comprising 70,000 grayscale images of size 28 × 28 pixels across 10 apparel categories (e.g., T-shirt/top, sneaker) with a 60,000/1,000 train/test split.
SVHN contains approximately 600,000 real-world 32 × 32 RGB digit crops (0–9) collected from Google Street View images. It is partitioned into training (73,257), testing (26,032), and an extra set of 531131 samples for data augmentation.
CIFAR-10 consists of 60,000 32 × 32 color images evenly distributed over 10 object classes (airplane, car, bird, cat, deer, dog, frog, horse, ship, truck). There are five training batches of 10000 images each and one test batch, with exactly 1000 images per class.
CIFAR-100 (building on CIFAR-10) contains 60,000 32 × 32 color images in 100 fine classes (600 images each) grouped into 20 coarse superclasses. Each fine class has a 500/100 train/test split, enabling hierarchical and fine-grained classification studies.
Tiny ImageNet is a subset of the ILSVRC-2012 challenge, selecting 200 classes and resizing all images to 64 × 64 pixels. It provides 100,000 images (500 train, 50 val, 50 test per class), offering a mid-scale benchmark between CIFAR and full ImageNet.
ImageWoof is a challenging subset of 10 visually similar dog breeds drawn from ImageNet (e.g., Beagle, Samoyed, Golden Retriever). It contains 9,025 training and 3,929 validation images, with optional noisy-label variants, and is commonly used to benchmark fine-grained recognition models.
@@ Appendix I Results on TinyImageNet
Random [7] 1.4$\pm$0.1 5.0$\pm$0.2 15.0$\pm$0.4
Herding [58] 2.8$\pm$0.2 6.3$\pm$0.2 16.7$\pm$0.3
K-Center [49] 1.6$\pm$0.2 5.1$\pm$0.1 15.0$\pm$0.3
Forgetting [55] 1.6$\pm$0.2 5.1$\pm$0.3 15.0$\pm$0.1
DC [69] 5.3$\pm$0.1 12.9$\pm$0.1 12.7$\pm$0.4
DSA [67] 5.7$\pm$0.1 16.3$\pm$0.2 5.1$\pm$0.2
DataDAM [48] 8.3$\pm$0.4 18.7$\pm$0.3 28.7$\pm$0.3
MTT [4] 6.2$\pm$0.4 17.3$\pm$0.2 26.5$\pm$0.3
IDM [70] 10.1$\pm$0.2 21.9$\pm$0.6 26.9$\pm$0.2
IDM with HDD 11.9$\pm$0.2 22.4$\pm$0.3 27.8$\pm$0.3
Whole Dataset 37.6$\pm$0.6
@@
Appendix J Visualization of Distilled Images

### v2

@@ A.1 Tangent Space $T_{\mathbf{x}}\mathcal{L}$
In the Lorentz model, hyperbolic space $\mathcal{L}$ is realized as a sheet of the two‐sheeted hyperboloid in $\mathbb{R}^{n+1}$ with Minkowski metric. For any point $\mathbf{x}=[x_{t};x_{s}]\in\mathcal{L}$, the tangent space captures all possible instantaneous directions at $\mathbf{x}$. It is defined by
@@ A.2 Exponential and Logarithm Maps
The exponential map pushes vectors in the tangent space onto the manifold, yielding a local Euclidean‐like parametrization. Let $\kappa=\sqrt{-K}$. For $\mathbf{v}\in T_{\mathbf{x}}\mathcal{L}$, define
@@ A.3 Bijection between the Lorentz and Poincaré Ball Models
For many applications (especially visualization), it is convenient to switch to the Poincaré ball. Given a Lorentz point $\mathbf{x}=[x_{t};x_{s}]$, we map it to the unit ball $\|\mathbf{p}\|<1$ via
@@
One verifies that the reconstructed $\mathbf{x}$ satisfies $x_{t}^{2}+\|x_{s}\|^{2}=1/K$ and $x_{t}>0$.
@@ Appendix B Centroid Convergence Toward the Origin
forcing the solution of $\sum_{i}\ Log_{\mathbf{p}}(\mathbf{p}_{i})=0$ to lie radially closer to $p_{0}$ than the Euclidean centroid. Moreover, as $\lvert K\rvert$ increases, the lower bound on the second‐derivative of $t\mapsto\|\ Log_{\gamma(t)}(\mathbf{p}_{i})\|^{2}$ along any geodesic $\gamma$ grows, making this radial bias toward $p_{0}$ even more pronounced. This results in the centroids of both the original dataset and the synthetic dataset being biased towards $p_{0}$, while the distance between them is relatively small.
@@ Appendix C Hierarchical Weight
$\mathcal{L}:=\bigl\{x\in\mathbb{R}^{n+1}\mid\langle x,x\rangle_{\mathcal{L}}=-R^{2},\;x_{0}>0\bigr\},$ (31)
where $\langle\cdot,\cdot\rangle_{L}$ denotes the Minkowski inner product of signature $(-+\cdots+)$. The geodesic distance between two points $p,q\in\mathbb{H}_{K}^{n}$ is given by
@@ Appendix D Gradient Contributions in the Lorentz Model of Hyperbolic Space
$f(t)\;=\;\frac{\ arcosh(t)}{\sqrt{t^{2}-1}}\;\longrightarrow\;1.$ (44)
Thus, points very close to $x$ exert almost the maximal pull of magnitude 1.
@@ Appendix E Benchmark Datasets
FashionMNIST is a drop‑in replacement for the classic MNIST dataset, comprising 70,000 grayscale images of size 28 × 28 pixels across 10 apparel categories (e.g., T‑shirt/top, sneaker) with a 60,000/1,000 train/test split.
SVHN contains approximately 600,000 real‑world 32 × 32 RGB digit crops (0–9) collected from Google Street View images. It is partitioned into training (73,257), testing (26,032), and an extra set of 531131 samples for data augmentation.
CIFAR‑10 consists of 60,000 32 × 32 color images evenly distributed over 10 object classes (airplane, car, bird, cat, deer, dog, frog, horse, ship, truck). There are five training batches of 10000 images each and one test batch, with exactly 1000 images per class.
CIFAR‑100 (building on CIFAR‑10) contains 60,000 32 × 32 color images in 100 fine classes (600 images each) grouped into 20 coarse superclasses. Each fine class has a 500/100 train/test split, enabling hierarchical and fine‑grained classification studies.
Tiny ImageNet is a subset of the ILSVRC‑2012 challenge, selecting 200 classes and resizing all images to 64 × 64 pixels. It provides 100,000 images (500 train, 50 val, 50 test per class), offering a mid‑scale benchmark between CIFAR and full ImageNet.
ImageWoof is a challenging subset of 10 visually similar dog breeds drawn from ImageNet (e.g., Beagle, Samoyed, Golden Retriever). It contains 9,025 training and 3,929 validation images, with optional noisy‑label variants, and is commonly used to benchmark fine‑grained recognition models.
@@ Appendix I Results on TinyImageNet
[…]
@@
Appendix J Experiments on an Alternative Hybrid Architecture (DSDM)
We evaluated the performance of HDD on the hybrid architecture DSDM [35] using the CIFAR-10 dataset. As shown in Table 10, when IPC = 1, DSDM with HDD achieved a 2.6% performance gain compared to the original DSDM; when IPC = 10, DSDM with HDD showed an improvement of 0.8%.
Table 10: Accuracy comparison of DSDM with/without HDD.
Method IPC Accuracy (%)
DSDM 1 43.8 $\pm$ 0.2
DSDM with HDD 1 46.4 $\pm$ 0.3
DSDM 10 65.8 $\pm$ 0.3
DSDM with HDD 10 66.6 $\pm$ 0.4
DSDM 50 75.8 $\pm$ 0.2
DSDM with HDD 50 76.0 $\pm$ 0.2
Appendix K Visualization of Distilled Images